OpenStudy (anonymous):

Winchesters Community Bank charges its customers $6 per month plus 15 cents per check for the use of a standard checking account. Thomas had a checking accound there for 8 months and was charged $53.10 in service charges. How many checks did he write during that period?

OpenStudy (anonymous):

8 month = 8*6=48 53.10-48=5.1 5.1/0.15=35 He write 35 checks